Train A is 30 miles ahead of Train B. Train A is traveling at 100 mph and train B is traveling at 120 mph. How long will it take Train B to overtake Train A?
Answer by mananth(16949)

Train A= 100 mph
Train B = 120 mph
catch up speed = 120-100 = 20 mph
Train A is ahead by 30 miles
Catch up time = catchup distance / catch up speed
time = 30/20
=1.5 hours